OpenStudy (crashonce):

The height and width of a cuboid are each increased by 10% and its length is decreased by 10%. The percentage increase in volume is?

OpenStudy (crashonce):

@perl

OpenStudy (danjs):

Volume of Cuboid = L*W*H right?

OpenStudy (crashonce):

yes

OpenStudy (danjs):

Height increases -- (H + 0.1H) Width Increases --- (W + 0.1W) Length Decreases --- (L - 0.1L)

OpenStudy (danjs):

Multiply those values 100% Volume = L x W x H x% volume = ....

OpenStudy (danjs):

Modified Volume = (1.089)(L x W x H)

OpenStudy (danjs):

Modified Volume / 100% Volume =

OpenStudy (danjs):

You could also try using numbers,, let original L, W, and H = 10, and the modified versions H = 11 W = 11 L = 9 Calculate and find the percent diff

OpenStudy (danjs):

\[\frac{ Modified }{ Original } = \frac{ 11 * 11 * 9 }{ 10^3 } = \frac{ 1189 }{ 1000 } \approx 1.089\]

OpenStudy (danjs):

modified is 108.9 percent the original volume

OpenStudy (danjs):

sound good?

OpenStudy (crashonce):

yes thanks